Varriano Corporation bases its budgets on the activity measure customers served. During October, the company planned to serve 48,000 customers, but actually served 49,000 customers. The company has provided the following data concerning the formulas it uses in its budgeting:

Fixed element
per month
Variable element per customer
Revenue - $ 3.86
Wages and salaries $ 38,200 $ 1.33
Supplies $ 0 $ 0.63
Insurance $ 8,700 $ 0.00
Miscellaneous expense $ 8,200 $ 0.53

Prepare a report showing the company's activity variances for October. Indicate in each case whether the variance is favorable (F) or unfavorable (U). (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

Varriano Corporation
Activity Variances
For the Month Ended October 31
Flexible Budget Planning Budget Activity Variances
Customers served 49,000 48,000
Revenue
Expenses:
Wages and salaries
Supplies
Insurance
Miscellaneous expense
Total expense
Net operating income
Answer #1

Answer-

VARRIANO CORPORATION
Activity Variance
For the month ended October 31
Particulars Planning Budget Flexible Budget Variances Remark
$ $ $
Revenue 48000 customers served*$3.86 per customer= 185280 49000 customers served*$3.86 per customer= 189140 3860 Favorable
Less:-Expenses
Wages & salaries $38200+(48000 customers served*$1.33 per customer)= 102040 $38200+(49000 customers served*$1.33 per customer)= 103370 -1330 Unfavorable
Supplies 48000 customers served*$0.63 per customer= 30240 49000 customers served*$0.63 per customer= 30870 -630 Unfavorable
Insurance 8700 8700 0 None
Miscellaneous expense $8200+(48000 customers served*$0.53 per customer)= 33640 $8200+(49000 customers served*$0.53 per customer)= 34170 -530 Unfavorable
Total Expenses 174620 177110 -2490 Unfavorable
Income from operations 10660 12030 1370 Favorable